# Quantum chaos in $\mathcal{PT}$-symmetric quantum systems

### v2

@@ 1 Introduction
Over the years, the notion of chaos in a quantum system has been placed on a firm footing despite the absence of a phase space in which to describe the dynamics,[1, 2]. Diagnostics such as the energy level spacing distribution have been employed to characterize chaotic quantum Hamiltonians due to their similarities with random matrices [3, 4]. Random Matrix Theory (RMT) is a powerful tool that accurately describes the spectral statistics of quantum systems whose classical counterparts exhibit chaotic behaviour. In cases where quantum Hamiltonians correspond to integrable classical systems, the Berry-Tabor conjecture proposes that their level spacing distributions have Poissonian forms,[5]. On the other hand, for quantum Hamiltonians with chaotic classical counterparts, the Bohigas-Giannoni-Schmit conjecture proposes that the level statistics should correspond to one of the three classical ensembles of RMT (Wigner-Dyson ensembles) [6], namely – the Gaussian unitary ensemble (GUE), the Gaussian orthogonal ensemble (GOE), and the Gaussian symplectic ensemble (GSE) that come out of Wigner’s surmise,[7, 8]. A similar diagnostic is the energy level spacing ratio, which has proven more versatile due to its independence from local energy densities,[9].
@@
Non-Hermitian Hamiltonians have been studied extensively over the last few years due to the discovery of phenomena such as the non-Hermitian skin effect [43], the presence of exceptional points [44] and interesting topological properties in these systems. [45, 46]. Bender showed that if a Hamiltonian commutes with the operator $\mathcal{PT}$ where $\mathcal{P}$ is the (unitary) parity operator and $\mathcal{T}$ is the (anti-unitary) time reversal operator, it can possess a completely real spectrum of eigenvalues without necessarily being Hermitian [47, 48]. In fact, one can typically choose a parameter to tune in such a $\mathcal{PT}$ symmetric Hamiltonian to go from a $\mathcal{PT}$ symmetric phase to a $\mathcal{PT}$ symmetry broken phase, in which the energy eigenvalues are complex [47].
It is thus interesting to investigate whether a system can exhibit both an integrable to chaotic transition and a $\mathcal{PT}$ symmetry-breaking transition, and if so, what are the possible resultant phases? Motivated by the above question, in this work, we study a non-Hermitian extension of the kicked rotor model, which we henceforth refer to as the $\mathcal{PT}$ symmetric kicked rotor (PTKR) model.
Since the PTKR model is non-Hermitian and may have complex eigenvalues, we must use generalizations of the previously discussed diagnostics of quantum chaos [49] that apply to complex eigenvalues. For example, we study the complex energy level spacing ratio (CLSR) [50, 51] instead of the conventional real energy level spacing ratio (RLSR) [9]. Additionally, we calculate a suitably defined OTOC for non-Hermitian systems. [52]. Recently, the time evolution of the OTOC has been studied for the PTKR model and its classical counterpart in the vicinity of a $\mathcal{PT}$ symmetry-breaking transition in parameter space.[53, 52, 54]. Our study extends this study to include the entire phase diagram of the PTKR model. It provides critical insights into the interplay of the $\mathcal{PT}$ symmetry-breaking transition and the transition from integrability to chaos.
@@ 2 Model
We modify the Hermitian quantum kicked rotor model by adding a non-Hermitian term that preserves $\mathcal{PT}$ symmetry. The resultant model is the PTKR model described by the Hamiltonian
@@
Here $\theta$ is the angle the rotor makes with a pre-defined direction and the momentum $p=-i\hbar d/d\theta$. When set to zero, the non-Hermiticity parameter $\lambda$ gives the Hermitian model. The action of the parity operator $\mathcal{P}$ is defined as follows-
@@
where we have taken $\hbar=1,\tau=1$ and $m=1$. The above-defined operator is time-independent and can be diagonalized to obtain its eigenvalues. To extract the eigenvalues of the PTKR model Hamiltonian from its Floquet evolution operator, we take the logarithm of the operator’s eigenvalues and multiply them by $i$. Due to the periodic nature of phases, the eigenvalue spectrum of the Floquet Hamiltonian is folded into a single interval modulo the period $2\pi$. Since this procedure causes eigenvalues that would otherwise have been far apart to now be proximate, it generates chaos.
@@ 3.1 Energy Level Spacing Ratio
To extend the analysis of the energy level spacing distribution to non-Hermitian systems, we use the complex level spacing ratio $\xi$ [50] defined as
$\xi_{K}=\dfrac{z_{K}^{\text{NN}}-z_{K}}{z_{K}^{\text{NNN}}-z_{K}}=r_{K}\exp{i\theta_{K}}$ ) (7)
The complex level spacing ratio (CLSR) can be defined in two ways: One is by averaging over the magnitudes of $\xi$ to obtain $\langle r\rangle$ (derived from the set of $r_{K}$ values), and the other is averaging over the angular distribution to obtain $-\langle\cos{\theta}\rangle$ ) ⟩ (derived from the set of $\theta_{K}$ values). Only $\langle r\rangle$ can be defined in the limit when the spectrum becomes real. However, in this limit, Eq. 7 does not yield the standard real level spacing ratio (RLSR) for a real spectrum. We, henceforth, refer only to the mean value $\langle r\rangle$ as the complex level spacing ratio (CLSR). Calculations of $-\langle\cos{\theta}\rangle$ ) ⟩ can be found in the supplementary material. The universality classes for random matrices that describe Hermitian matrices in terms of Gaussian Ensembles have suitable generalization to complex matrices [49]. We calculate the CLSR for these universality classes, one of which the PTKR model belongs.
There are two computational techniques to improve the statistics in our calculation. The first is to replace $m$ in Eq. (1) by $m+\Delta m_{p}$, where $\Delta m_{p}$ is a small random number selected independently for each $p$. This ensures that no underlying symmetries leading to unwanted degeneracies occur. This is important as degeneracies must be avoided when calculating level spacing ratios. The second is to average over a small window of values of the kicking strength $K$. This is done to avoid any accidental resonance in the $K$ values. Although, this averaging wasn’t performed while plotting the colormaps to keep the transitions intact.
@@ 3.2 Out-of-time-order correlator
Figure 1: Left The CLSR and RLSR as a function of the kicking strength $K$, calculated for $N=8005$ with $\hbar=0.2$. Horizontal lines represent analytical values of RLSR for Poisson and GOE statistics. The figure shows that the CLSR is as good an indicator of the transition from integrability to chaos as the more commonly employed RLSR. Right The CLSR and RLSR as a function of the non-Hermiticity parameter $\lambda$, calculated for $N=4095$ with $\hbar=0.2$ and $K=0.15$. Horizontal lines represent the values of CLSR for the Poisson and GOE distributions. The figure resembles the transition observed for the CLSR on the left got by varying $K$ with $\lambda=0$. However, it is obtained by tuning $\lambda$ instead of $K$. This suggests that the system can transition from a $\mathcal{PT}$-symmetric integrable phase to a $\mathcal{PT}$-symmetric chaotic phase with increasing non-Hermiticity while the value of $K$ is kept constant.
@@
Figure 2: Left: The CLSR for varying values of the kicking strength $K$ and the non-Hermiticity $\lambda$, computed for $N=6005$, $\hbar=0.2$. The dotted lines T1 and T2 represents the transitions across the $\mathcal{PT}$-symmetric integrable phase to $\mathcal{PT}$ symmetry broken chaotic phase and the $\mathcal{PT}$-symmetric chaotic phase to $\mathcal{PT}$ symmetry broken chaotic phase, respectively. Line T* is the Hermitian transition from the integrable to chaotic phase. T3 marks a transition similar to T* in which $K$ is held constant and $\lambda$ varied. Center: The CLSR across the transition Main panel: T1 Inset T2, calculated for $N=4095$. The horizontal lines show the standard values of the CLSR for the GOE and Poisson distributions. Right: The absolute value of the maximum imaginary part of an energy eigenvalue, $\alpha$, across the transition Main panel: T1 Inset T2, calculated for $N=4095$. It can be seen that while $\alpha$ shows an abrupt change along T2, it seems to increase smoothly along T1. The CLSR on the other hand, shows an abrupt transition along both T1 and T2.
@@ 4.1 CLSR
In Fig. 1, we show the mean RLSR as well as the CLSR with varying values of the kicking strength $K$, for $\hbar=0.2$ and system-size $N=8005$. We notice that the transition points are reasonably independent of the value of $N$. Thus, in what follows, the system size is chosen to be large enough so that all quantities are well converged. We find that both the CLSR and RLSR display a transition at the same value of $K$. However, the values in the two regimes are different. Furthermore, the values of the CLSR and RLSR that are obtained by the PTKR model in the chaotic limit of the Hermitian case match with those obtained from averaging over purely random matrices in the Gaussian Orthogonal Ensemble (GOE). The transition in the Hermitian case has been highlighted in the color plot as T* for aid in comparison.
Table 1: The above table consists of data gathered on CLSR and RLSR of standard known values in the integrable regime and random matrix ensemble calculations in the chaotic regime [9, 51].
@@
It is interesting to note that the transition from the $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetric chaotic also occurs across $T_{3}$. The hermitian case $\lambda=0$ exhibits a transition from $\mathcal{PT}$- symmetric integrable to the $\mathcal{PT}$-symmetric chaotic phase as indicated by $T*$. A similar transition can thus be expected when one varies $K$ for $\lambda$ sufficiently small but not equal to zero as seen in the left panel of Fig. 2. The same transition can also be effected by varying $\lambda$ but not $K$ as indicated by $T_{3}$. This is somewhat surprising as it seems to lack any Hermitian counterpart. One is thus able to induce chaos, as seen from the measured CLSR values, by increasing $\lambda$ whilst holding $K$ to the traditional integrable value. The nature of this transition is intriguing and requires further study, which we defer to future work.
Figure 3: The OTOC $C(t)$ and state norm $\mathcal{N}(t)$ vs $t$ across the transition along Left $T_{1}$ and Right $T_{2}$ shown in Fig. 2, calculated for system size $N=2^{14}$. For T1 (Left), the values $(\lambda,K)=\{(0.001,0.1),(0.002,0.456),(0.003,1.135)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken integrable phase, while the values $(\lambda,K)=\{(0.006,5.734),(0.01,15.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. For T2 (right), the values $(\lambda,K)=\{(10^{-5},30),(10^{-4},30),(10^{-3},30)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken chaotic phase, while the values $(\lambda,K)=\{(0.005,30),(0.001,30.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. In both cases, the $\mathcal{PT}$-symmetric broken phase displays a late-time exponential growth of the OTOC and the state norm.
@@ 4.2 OTOC
We now present our results from the calculation of the OTOC. More specifically, we present the OTOC calculations for various phases obtained from the CLSR analysis and across the transitions $T_{1}$ and $T_{2}$.
@@
$\mathcal{PT}$-symmetric integrable phase: The Hermitian case $\lambda=0$ has been studied previously  [42], where it was shown that the OTOC follows a power law for all time scales in the integrable phase. On the other hand, the OTOC features a transition from exponential growth at early times $t<t_{E}$ to a power law in the chaotic phase. In Fig. 3 (right), we show the OTOC $C(t)$ for the $\mathcal{PT}$-symmetric integrable phase with parameters $(\lambda,K)\in\{(0.001,0.1),(0.002,0.456),\allowbreak(0.03,1.135)\}$, and system-size $2^{14}$. We find that the OTOC exhibits a behavior similar to the Hermitian case, which agrees with the calculations of the CLSR.
@@
$\mathcal{PT}$-symmetric chaotic phase: Figure 3 (right) shows the OTOC for the $\mathcal{PT}$-symmetric chaotic phase for $\lambda=(10^{-5},10^{-4})$, and fixed $K=30$. As for the $\mathcal{PT}$-symmetric integrable phase, we find that the OTOC in the symmetric chaotic phase exhibits behavior similar to that of the Hermitian system in agreement with CLSR calculations.
$\mathcal{PT}$-symmetry broken chaotic phase: This is the only $\mathcal{PT}$-symmetry broken phase present in the phase diagram. In Fig. 3 (left), we show the OTOC for $\lambda=(10^{-3},0.005,0.001)$, and fixed kicking strength $K=30$. In Fig. 3 (right), the OTOC for $(\lambda,K)\in\{(0.006,5.734),(0.01,15)\}$ is shown. It displays two separate regimes of exponential growth. In early times, the growth is dominated by the chaotic nature of the Hamiltonian. In contrast, at late times, the growth is essentially because of the complexity of eigenvalues brought about by $\mathcal{PT}$-symmetry breaking, as we explain below.
In the $\mathcal{PT}$-symmetry broken phase, the eigenspectrum consists of complex eigenvalues, and thus under time evolution, the state $|\Psi\rangle$ no longer remains normalized (See Fig. 3). The late time growth of the OTOC (and state norm) depends on the eigenspectrum, in particular, on the maximum of the imaginary part of the eigenvalues of the Hamiltonian. Let us denote this maximum value by $\alpha$. Consider the late-time behavior of the state $|\Psi\rangle$
Since, at late times, the evolution is dominated by the phase $e^{-iEt}$ from $E=E_{r}+i\alpha$, it follows that the late time norm $\mathcal{N}(t)$ obeys
Figure 4: The normalized OTOC $\tilde{C}(t)$ vs $t$ corresponding to the OTOC $C(t)$ shown in Fig. 3. At late times, $\tilde{C}(t)$ has a constant value after accounting for the exponential growth resulting from broken $\mathcal{PT}$-symmetry.
$\mathcal{N}(t\rightarrow\infty)\propto e^{2\alpha t}\,.$ (11)
To consider the behavior of the OTOC, we explicitly write the commutation relation $-[\hat{p}(t),\hat{p}(0)]^{2}$ as
In the large $t$ limit, we can write the contribution from a single term as
Therefore, the late time OTOC
$C(t\rightarrow\infty)\propto e^{4\alpha t}\,$ (14)
This behavior motivates a definition of a normalized OTOC
$\tilde{C}(t)=e^{-4\alpha t}C(t)\,.$ (15)
@@
The normalized OTOC, $\tilde{C}(t)$, is expected to exhibit exponential growth only due to the chaotic nature of the Hamiltonian. Since we are only performing calculations up to finite times, there is a contribution from sub-leading eigenvalues as well. Therefore, the exponent is obtained from an exponential fitting only for late times (see supplementary material). The normalized OTOC, then, only contains a growth due to the chaotic nature of the system (See Fig. 4) and can be used to extract a Lyapunov exponent. Numerically, in order to extract the exponent from $\tilde{C}(t)$, we determine the times, after which the exponential growth starts slowing down, and fit $\tilde{C}(t)$ from $t=1$ up to these times to the function $ae^{2\gamma(t-1)}$ to find the parameter $\gamma$. Numerical overflows prevent a calculation of the Lyapunov exponent if it has a large value. We, thus, have to confine ourselves to a calculation of the exponent only in a region of the phase where $\mathcal{PT}$-symmetry is broken weakly.
@@
In Figure 6, we show the Lyapunov exponent as well as the corresponding CLSR across the transition $T_{2}$ for varying $\hbar_{\text{eff}}$. We find that for all values of $\hbar_{\text{eff}}$, the Lyapunov exponent displays a peak. It is important to note that the peak does not appear at the transition $T_{2}$ but instead exists inside the broken phase of the $\mathcal{PT}$ symmetry. A further investigation of the nature of the peak in the Lyapunov exponent profile (See Sec. 5) is required. We also note that in a calculation performed on the same system as ours [55], a different definition of the OTOC has been used, one involving both the operators $\theta$ and $p$. The OTOC defined this way displays polynomial growth in time in the $\mathcal{PT}$ symmetry broken regime and not an exponential one, as we observe.
Figure 5: The above figure shows a section of the CLSR contour map for different values of $\hbar_{\text{eff}}$. It can be observed that the transition T* moves closer to the integrable $\to$ chaotic transition seen in the classical model for which $K\sim 1$ as $\hbar_{\text{eff}}\to 0$. Furthermore, the PT-symmetric chaotic regime is observed to shrink in size with a decrease in $\hbar_{\text{eff}}$

### v3

@@ 1 Introduction
Over the years, the notion of chaos in a quantum system has been placed on a firm footing despite the absence of a phase space in which to describe the dynamics,[1, 2]. Diagnostics such as the energy level spacing distribution have been employed to characterize chaotic quantum Hamiltonians due to their similarities with random matrices [3, 4]. Random Matrix Theory (RMT) is a powerful tool that accurately describes the spectral statistics of quantum systems whose classical counterparts exhibit chaotic behaviour. In cases where quantum Hamiltonians correspond to integrable classical systems, the Berry-Tabor conjecture proposes that their level spacing distributions have Poissonian forms,[5]. On the other hand, for quantum Hamiltonians with chaotic classical counterparts, the Bohigas-Giannoni-Schmit conjecture proposes that the level statistics should correspond to one of the three classical ensembles of RMT (Wigner-Dyson ensembles) [6], namely – the Gaussian unitary ensemble (GUE), the Gaussian orthogonal ensemble (GOE), and the Gaussian symplectic ensemble (GSE) that come out of Wigner’s surmise,[7, 8]. Wigner’s surmise is a statement about the probability density of nearest-neighbor energy level spacings of Hermitian random matrices. Based on the symmetries of the matrix, the probability density takes one of three forms which can then be used to deduce information about the symmetries from the energy level spacing distribution. A similar diagnostic is the energy level spacing ratio, which has proven more versatile due to its independence from local energy densities[9].
@@
Non-Hermitian Hamiltonians have been studied extensively over the last few years due to the discovery of phenomena such as the non-Hermitian skin effect [43], the presence of exceptional points [44] and interesting topological properties in these systems  [45, 46]. The study of the effect of electron-electron interactions in non-Hermitian systems with Lorentz symmetry [47, 48, 49, 50] also proves to be of interest as it leaves the eigenspectrum of the system completely real. Bender showed that if a Hamiltonian commutes with the operator $\mathcal{PT}$ where $\mathcal{P}$ is the (unitary) parity operator and $\mathcal{T}$ is the (anti-unitary) time reversal operator, it can possess a completely real spectrum of eigenvalues without necessarily being Hermitian [51, 52]. In fact, one typically chooses a parameter to tune in such a $\mathcal{PT}$-symmetric Hamiltonian to go from a $\mathcal{PT}$-symmetric phase to a $\mathcal{PT}$-symmetry broken phase, in which the energy eigenvalues are complex [51].
It is thus interesting to investigate whether a system can exhibit both an integrable to chaotic transition and a $\mathcal{PT}$-symmetry-breaking transition, and if so, what are the possible resultant phases? Motivated by the above question, in this work, we study a non-Hermitian extension of the kicked rotor model, which we henceforth refer to as the $\mathcal{PT}$-symmetric kicked rotor (PTKR) model.
Since the PTKR model is non-Hermitian and may have complex eigenvalues, we must use generalizations of the previously discussed diagnostics of quantum chaos [53] that apply to complex eigenvalues. For example, we study the complex energy level spacing ratio (CLSR) [54, 55] instead of the conventional real energy level spacing ratio (RLSR) [9]. Additionally, we calculate a suitably defined OTOC for non-Hermitian systems. [56]. Recently, the time evolution of the OTOC has been studied for the PTKR model and its classical counterpart in the vicinity of a $\mathcal{PT}$ symmetry-breaking transition in parameter space.[57, 56, 58].Our work has also been inspired by previous work on the Hermitian-kicked rotor [42], which presents a numerical calculation of the Out of Time Ordered Correlator (OTOC) to the study the transition from integrability to chaos in the model. Our study extends this study to include the entire phase diagram of the PTKR model. It provides critical insights into the interplay of the $\mathcal{PT}$-symmetry-breaking transition and the transition from integrability to chaos. We study the interplay of these transitions by calculating the Complex level Spacing Ratio (CLSR) introduced in [54]. Additionally, we study a modified version of the OTOC by taking into account the lack of unitarity in the non-Hermitian systems to show possible markers of these transitions that can be captured by the OTOC.
@@ 2 Model
We modify the Hermitian quantum kicked rotor model by adding a non-Hermitian term that preserves $\mathcal{PT}$-symmetry. The resultant model is the PTKR model described by the Hamiltonian
@@
Here $\theta$ is the angle the rotor makes with a pre-defined direction and the momentum $p=-i\hbar_{\text{eff}}d/\allowbreak d\theta$. When set to zero, the non-Hermiticity parameter $\lambda$ gives the Hermitian model. The action of the parity operator $\mathcal{P}$ is defined as follows-
@@
where we have taken $\hbar_{\text{eff}}=1,\tau=1$ and $m=1$. The above-defined operator is time-independent and can be diagonalized to obtain its eigenvalues. To extract the eigenvalues of the PTKR model Hamiltonian from its Floquet evolution operator, we take the logarithm of the operator’s eigenvalues and multiply them by $i$. Due to the periodic nature of phases, the eigenvalue spectrum of the Floquet Hamiltonian is folded into a single interval modulo the period $2\pi$. Since this procedure causes eigenvalues that would otherwise have been far apart to now be proximate, it generates chaos.
@@ 3.1 Energy Level Spacing Ratio
To extend the analysis of the energy level spacing distribution to non-Hermitian systems, we use the complex level spacing ratio $\xi$ for the $\gamma^{\text{th}}$ state [54] defined below where $z_{\gamma}$ is its complex eigenvalue.
$)\xi_{\gamma}=\dfrac{z_{\gamma}^{\text{NN}}-z_{\gamma}}{z_{\gamma}^{\text{NNN}}-z_{\gamma}}=r_{\gamma}\exp{i\theta_{\gamma}}$ (7)
The complex level spacing ratio (CLSR) can be defined in two ways: One, by averaging over $\gamma$ the magnitudes of $\xi_{\gamma}$ to obtain $\langle r\rangle$ (derived from the set of $r_{\gamma}$ values), and the other, by averaging over the angular distribution to obtain $)-\langle\cos{\theta}\rangle$ (derived from the set of $\theta_{\gamma}$ values). Only $\langle r\rangle$ can be defined in the limit when the spectrum becomes real. However, in this limit, Eq. 7 does not yield the standard real level spacing ratio (RLSR) for a real spectrum. We, henceforth, refer only to the mean value $\langle r\rangle$ as the complex level spacing ratio (CLSR). Calculations of $)-\langle\cos{\theta}\rangle$ can be found in the supplementary material. The universality classes for random matrices that describe Hermitian matrices in terms of Gaussian ensembles have suitable generalization to complex matrices [53]. We calculate the CLSR for these universality classes, one of which the PTKR model belongs.
To improve the statistics in our calculation, we replace $m$ in Eq. (1) by $m+\Delta m_{p}$, where $\Delta m_{p}$ is a small random number selected independently for each momentum eigenstate |$\ket{p}$. This ensures that no underlying symmetries leading to unwanted degeneracies occur. This is important as degeneracies must be avoided when calculating level spacing ratios.
@@ 3.2 Out-of-time-order correlator
Figure 1: Left The CLSR and RLSR as functions of the kicking strength $K$, calculated for $N=10095$ with $\hbar_{\text{eff}}=0.2$ and $\lambda=0$ (Hermitian). Horizontal lines in orange (blue) correspond to values of the RLSR (CLSR) for Poisson and GOE statistics. The figure shows that the CLSR is as good an indicator of the transition from integrability to chaos as the more commonly employed RLSR. Right The CLSR as a function of the non-Hermiticity parameter $\lambda$, calculated for $N=10095$ with $\hbar_{\text{eff}}=0.2$ and $K=0.15$. Horizontal lines represent the values of CLSR for the Poisson and GOE distributions. The figure resembles the transition observed for the CLSR on the left obtained by varying $K$ with $\lambda=0$. However, it is obtained by tuning only $\lambda$ instead of $K$. This suggests that the system can transition from a $\mathcal{PT}$-symmetric integrable phase to a $\mathcal{PT}$-symmetric chaotic phase with increasing non-Hermiticity while the value of $K$ is kept constant.
@@
Figure 2: (a) The CLSR for varying values of the kicking strength $K$ and the non-Hermiticity $\lambda$, computed for $N=6005$, $\hbar_{\text{eff}}=0.2$. The dashed lines $T_{2}$ and $T_{1}$ represent the transitions across the $\mathcal{PT}$-symmetric integrable phase to $\mathcal{PT}$-symmetry broken chaotic phase and the $\mathcal{PT}$-symmetric chaotic phase to $\mathcal{PT}$-symmetry broken chaotic phase, respectively. Line $T^{*}$ is the Hermitian transition from the integrable to chaotic phase. $T_{3}$ marks a transition similar to $T^{*}$ in which $K$ is held constant and $\lambda$ varied. (b): The CLSR across the transition Main panel: $T_{2}$ Inset $T_{1}$, calculated for $N=6005$. The horizontal lines show the standard values of the CLSR for the GOE and Poisson distributions. (c): The maximum imaginary part of an energy eigenvalue, $\alpha$, across the transition Main panel: $T_{2}$ Inset $T_{1}$, calculated for $N=6005$. It can be seen that while $\alpha$ shows an abrupt change along $T_{2}$, it seems to increase smoothly along $T_{1}$. The CLSR on the other hand, shows an abrupt transition along both $T_{2}$ and $T_{1}$.
@@ 4.1 CLSR
In Fig. 1, we show the mean RLSR as well as the CLSR with varying values of the kicking strength $K$, for $\hbar_{\text{eff}}=0.2$ and system-size $N=10095$. We notice that the transition points are reasonably independent of the value of $N$. Thus, in what follows, the system size is chosen to be large enough so that all quantities are well converged. We find that both the CLSR and RLSR display a transition at the same value of $K$. However, the values in the two regimes are different. We can understand why the CLSR and the RLSR need not yield the same ratio when the spectrum is completely real in the following manner: Without loss of generality let us assume that corresponding to the energy $E_{i}$ we have $E_{i}-E_{i-1}<E_{i+1}-E_{i}$ and we define $\Delta{i,n}=|E_{i}-E_{i+n}|$. The RLSR is then defined for $E_{i}$ as $\Delta_{i,-1}/\Delta_{i,1}$. However, it is possible that $E_{i}-E_{i-2}<E_{i+1}-E_{i}$ in which case the CLSR as defined for $E_{i}$ is $\Delta_{i,-1}/\Delta_{i,-2}$ since it involves the ratio of energy differences between an eigenvalue and its nearest and next nearest neighbors. We can thus clearly see that when averaged over all the energy eigenvalues, the CLSR is guaranteed to be greater than or at least equal to the RLSR. Furthermore, the values of the CLSR and RLSR that are obtained for the PTKR model in the chaotic limit of the Hermitian case match with those obtained from averaging over purely random matrices in the Gaussian Orthogonal Ensemble (GOE). The transition in the Hermitian case has been highlighted in the color plot as $T^{*}$ for aid in comparison.
Table 1: The table below consists of data gathered on CLSR and RLSR of standard known values in the integrable regime and random matrix ensemble calculations in the chaotic regime [9, 55].
@@
It is interesting to note that the transition from the $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetric chaotic also occurs across $T_{3}$. The hermitian case $\lambda=0$ exhibits a transition from $\mathcal{PT}$- symmetric integrable to the $\mathcal{PT}$-symmetric chaotic phase as indicated by $T^{*}$. A similar transition can thus be expected when one varies $K$ for $\lambda$ sufficiently small but not equal to zero as seen in the left panel of Fig. 2. The same transition can also be effected by varying $\lambda$ but not $K$ as indicated by $T_{3}$. This is somewhat surprising as it seems to lack any Hermitian counterpart. One is thus able to induce chaos, as seen from the measured CLSR values, by increasing $\lambda$ whilst holding $K$ to the traditional integrable value. The nature of this transition is intriguing and requires further study, which we defer to future work.
@@ 4.2 OTOC
We present our results for the OTOC, focusing on its behavior across the different phases identified via the CLSR analysis. In the $\mathcal{PT}$-symmetry broken phase, the emergence of complex eigenvalues leads to unbounded growth of the state norm under time evolution. To account for this, we define a normalized OTOC (see Appendix D), which removes the exponential growth arising purely from non-Hermiticity and isolates the contribution due to chaotic dynamics.
Figure 3: The normalized OTOC $\tilde{C}(t)$ vs $t$ across the transition along Top $T_{1}$ and Bottom $T_{2}$ shown in Fig. 2, calculated for system size $N=2^{14}$. For $T_{2}$ (bottom), the values $(\lambda,K)=\{(0.001,0.1),(0.002,0.456),(0.003,1.135)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken integrable phase, while the values $(\lambda,K)=\{(0.006,5.734),(0.01,15.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. For $T_{1}$ (top), the values $(\lambda,K)=\{(10^{-5},30),(10^{-4},30),(10^{-3},30)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken chaotic phase, while the values $(\lambda,K)=\{(0.005,30),(0.001,30.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase.
@@
$\mathcal{PT}$-symmetric integrable phase: The Hermitian case $\lambda=0$ has been studied previously  [42], where it was shown that the OTOC follows a power law for all time scales in the integrable phase. On the other hand, the OTOC features a transition from exponential growth at early times $t<t_{E}$ to a power law in the chaotic phase. In Fig. 3 (bottom), we show the normalized OTOC $\tilde{C}(t)$ for the $\mathcal{PT}$-symmetric integrable phase with parameters $(\lambda,K)\in\{(0.001,0.1),\allowbreak(0.002,0.456),\allowbreak(0.03,1.135)\}$, and system-size $2^{14}$. We find that the normalized OTOC exhibits a behavior similar to the Hermitian case, which agrees with the calculations of the CLSR.
@@
$\mathcal{PT}$-symmetric chaotic phase: Figure 3 (top) shows the normalized OTOC for the $\mathcal{PT}$-symmetric chaotic phase for $\lambda=(10^{-5},10^{-4})$, and fixed $K=30$. As for the $\mathcal{PT}$-symmetric integrable phase, we find that the normalized OTOC in the symmetric chaotic phase exhibits behavior similar to that of the Hermitian system in agreement with CLSR calculations.
$\mathcal{PT}$-symmetric broken chaotic phase: This is the only $\mathcal{PT}$-symmetry broken phase present in the phase diagram. In Fig. 3 (top), we show the normalized OTOC for $\lambda=(10^{-3},0.005,0.001)$, and fixed kicking strength $K=30$. In Fig. 3 (bottom), the normalized OTOC for $(\lambda,\allowbreak K)\in\{(0.006,\allowbreak 5.734),\allowbreak(0.01,15)\}$ is shown. Since the normalized OTOC eliminates the growth due to non-Hermiticity, it exhibits behavior similar to that in the $\mathcal{PT}$-symmetric unbroken chaotic phase. Below we the study behavior of the growth rate by calculating the Lyapunov exponent across the $T_{2}$ transition.
Figure 4: The above figure provides a comparison between the observed $\mathcal{PT}$-symmetric chaotic $\to$ PT-symmetry broken chaotic regime transition probed by the CLSR and the accompanying LE values for the same set of parameters for $K$ and $\lambda$. These have been plotted for various values of $\hbar_{\text{eff}}$ that span three orders of magnitude. We observe that this transition observed by the CLSR is accompanied by a peak in the LE. Note that different values of $K$ have only been used to ensure the transition is properly captured, since, as previously stated, the $\mathcal{PT}$-symmetric chaotic regime shrinks as $\hbar_{\text{eff}}\to 0$.
@@
The normalized OTOC captures only the growth due to chaotic dynamics and thus allows for the extraction of a Lyapunov exponent (LE). Numerically, in order to extract the exponent from $\tilde{C}(t)$, we determine the times, after which the exponential growth starts slowing down, and fit $\tilde{C}(t)$ from $t=1$ up to these times to the function $ae^{2\Lambda(t-1)}$ to find the parameter $\Lambda$. Numerical overflows prevent a calculation of the LE if it has a large value. We, thus, have to confine ourselves to a calculation of the exponent only in a region of the phase where $\mathcal{PT}$-symmetry is broken weakly.
@@
In Figure 4, we show the LE as well as the corresponding CLSR across the transition $T_{1}$ for varying $\hbar_{\text{eff}}$. We find that for all values of $\hbar_{\text{eff}}$, the LE displays a peak. It is important to note that the peak does not appear at the transition $T_{1}$ but instead exists inside the broken phase of the $\mathcal{PT}$-symmetry. A further investigation of the nature of the peak in the Lyapunov exponent profile (See Sec. 5) is required. We also note that in a calculation performed on the same system as ours [59], a different definition of the OTOC has been used, one involving both the operators $\theta$ and $p$. The OTOC defined this way displays polynomial growth in time in the $\mathcal{PT}$-symmetry broken regime and not an exponential one, as we observe.
Figure 5: The figure above shows the transition from the PT-symmetric integrable regime to the PT-symmetric chaotic regime for different system sizes that increase by a factor of 222. $\hbar_{\text{eff}}=0.2$ and $K=30$ for all the plots. It can be seen that the transition, as obtained from the CLSR gets sharper as the system size ($N$) increases. There is a slight shift in the value of $\lambda$ at which the transition occurs and this shift appears to decrease with successively increasing values of $N$. The peak in the LE also appears to remain at a nearly fixed value of $\lambda$ as the system size is varied.